# Two-photon interference between mutually-detuned resonance fluorescence signals scattered off a semiconductor quantum dot

### v2

@@ Abstract
Radiative linewidth of a two-level emitter (TLE) ultimately determines the bandwidth it can offer for quantum information processing. However, no prior experiment has so far been performed to examine the effect of driving detuning on indistinguishability of photons scattered off a TLE, a parameter that is crucial for photonic quantum computing. Here, we perform post-selective two-photon interference experiments between mutually-detuned resonance fluorescence signals from an InAs quantum dot embedded in a micropillar cavity. Our results suggest that indistinguishability among photons scattered off a quantum dot is inherently insensitive to the driving laser’s detuning, as straightforwardly predicted by the resonance fluorescence model that systematically treats all scattered photons as spontaneous emission.
@@
Impinging upon a semi-transparent beam-splitter, two indistinguishable single photons will coalescence into either of the two output ports, and never exit separately. Referred to as Hong-Ou-Mandel (HOM) effect 1, this phenomenon is the foundation for quantum information protocols such as repeater-assisted quantum communication 2 and linear-optic quantum computing 3. Single photons can be generated through exciting an atom-like system 4, 5, 6, which by nature emits one photon at a time. In this regard, semiconductor quantum dots (QDs) are arguably the most attractive platform because they can be integrated into monolithic structures for cavity enhancement 5, 7, 8 and highly indistinguishable single photons 9, 10, 11.
@@
Resonant excitation into the lowest excited state allows a QD system to be simplified as a two-level emitter (TLE) 12. Under a weak coherent pump (Heitler regime), the resulting resonance fluorescence (RF) signal locks to the driving laser’s coherence 13, 14, 15, 16 while exhibiting single-photon characteristic. To a first approximation, the QD appears to behave like a passive and elastic scatterer, which led to an optimistic idea to circumvent environment-induced dephasing inherent in all solid-state systems and to generate single photons with indistinguishability “fundamentally” defined by the excitation laser 14. A question then arises. In the presence of an excitation detuning, will the single photons remain indistinguishable to their counterparts generated under strictly resonant excitation? An affirmative answer would be favorable for quantum information processing because it will then permit arbitrary modulation of the driving laser and the shaping of the RF beam accordingly without loss of photon indistinguishability. However, no experiment has so far been dedicated to explicitly examine this question.
@@
Fig. 1: Dual-color two-photon interference setup. a, High-resolution RF spectra measured with a Fabry-Pèrot interferometer (FPI); The dashed line represents the QD-micropillar reflectance spectrum measured under $\bar{n}=0.05$. b, The auto-correlation function $g^{(2)}(0)$ values measured under $\bar{n}=0.05$ using a Hanbury Brown-Twiss setup (HBT). The red dashed line shows the cavity reflectivity under a strong excitation of 465 nW, corresponding to $\bar{n}=139$. BS: beam splitter; SPD: single-photon detector.
@@
Here, we subject an InAs QD-micropillar device to dual-color excitation, i.e., from two different wavelengths, using the setup shown in Fig. 1, and evaluate indistinguishability of mutually-detuned RF signals using post-selective two-photon interference (TPI) 17. Two continuous-wave (CW) lasers (not shown) are carved into 0.84 MHz square waves before being temporally interleaved to excite the QD device. Symmetrical detuning is applied: one laser is detuned by $+\frac{\Delta}{2}$ while the other by $-\frac{\Delta}{2}$ from the QD resonance ($\nu_{0}$). The resulting RF tracks the lasers’ frequencies and consists of RF signal pulses of alternating colors. The pulses enter an asymmetric Mach-Zehnder interferometer (AMZI) of 595 ns delay that matches the laser carving frequency, so that the RF signals of different colors are temporally aligned when meeting at the AMZI’s exit beam splitter. The interference results are registered by a pair of superconducting nanowire single photon detectors (SPDs) of 59 ps temporal resolution.
The InAs QD sample was fabricated from a wafer grown by molecular beam epitaxy on a (001) semi-insulating GaAs substrate and embedded into a double distributed-Bragg-reflector (DBR) micropillar cavity with a diameter of 2.4 $\mu$m and a quality factor (Q) of 9350. A confocal microscope (not shown) collects the RF at the same polarization as the excitation, thanks to the low cavity reflection of the device 18, 16. The sample is cooled by a closed-cycle cryostat to 5.6 K to ensure the QD’s resonance with the micropillar cavity of 911.485 nm. Under pulsed resonant excitation, we characterize the QD to have a radiative lifetime ($T_{1}$) of $74$ ps, corresponding to a natural linewidth of $\gamma_{\parallel}:=\frac{1}{2\pi T_{1}}=2.15$ GHz. In all subsequent experiments, we denote the excitation flux upon the device surface using the relation $\bar{n}=P_{\mathrm{in}}T_{1}/h\nu$, where $P_{\mathrm{in}}$ corresponds to the CW incident optical power.
@@
Before the dual-color experiment, we characterize the QD under excitation by a single laser. The excitation flux is kept at $\bar{n}=0.05$, which is conservatively set low to operate in the Heitler regime. Fig. 1a shows a set of high-resolution RF spectra measured by a Fabri-Pèrot interferometer (FPI). At each detuning, the RF peak inherits the laser’s optical frequency and has a full-width at half-maximum width (FWHM) that is limited by the FPI resolution ($20$ MHz). The peak intensity decreases as the detuning widens, which is a result of the reduced cavity enhancement. At the measurement step-size of 0.5 GHz, the adjacent RF spectra are fully distinguished. Fig. 1b shows the corresponding auto-correlation function results, measured with a Hanbury Brown-Twiss (HBT) setup. Across the detuning range, $g^{(2)}(0)$ is consistently lower than 0.5, indicating the non-classical nature of each RF signal. The lowest $g^{(2)}(0)=0.026$ is obtained at 0-detuning. It gradually deteriorates as the detuning increases, and exceeds $0.10$ when the detuning reaches ±$\pm 2$ GHz. We attribute the deterioration in single-photon purity to the increased laser background, as evidenced by the cavity reflectivity (red dashed line), which was measured under strong laser excitation.
@@
We use the setup in Fig. 1 for dual-color excitation experiments. By rotating the half-wave plate ($\lambda/2$) in the long arm of the AMZI, the polarization alignment between the incoming photon beams can be made either parallel or orthogonal. In the latter case, photons are fully distinguishable. We measure the normalized second-order correlation functions under both polarization settings, from which we calculate the TPI visibility ($\mathcal{V}$) using 17
@@
As a control experiment, we set the two lasers to be nominally identical in wavelength, i.e., $\Delta=0$ GHz, and Fig. 2a shows the measured HOM traces (dots). Here, the coincidence probabilities are normalized to the average rate at long delays ($\tau>200$ ns). With orthogonal polarizations, the photons are fully distinguishable and therefore no interference takes place. Hence, the corresponding $g^{(2)}_{\perp}(\tau)$ trace displays a single, sharp dip with a depth of $\sim$0.5 and a width of about 100 ps, arising from the single-photon nature of the RF streams. When the polarizations are aligned, photons become largely indistinguishable, two-photon coalescence takes place thus resulting in a more pronounced HOM dip at $\tau=0$. Additionally, the correlation function exhibits a broad suppression extending over a timescale of ±$\pm 100$ ns. Over nanosecond timescale (Fig. 2c), $g_{\perp}^{(2)}$ and $g_{\parallel}^{(2)}$ exhibit flat values for $\tau\gg T_{1}$ but with $g_{\parallel}^{(2)}$ being considerably lower than $g_{\perp}^{(2)}$. This is a common observation in the HOM measurements for resonantly excited QDs 16, 15. Using Eq. 1, we obtain the TPI visibility shown in Fig. 2e.
@@
Next, we measure the second-order HOM correlation functions under dual-color excitation, with the result of a $\Delta=0.5$ GHz detuning shown in Fig. 2b. While the $g^{(2)}_{\perp}$ data is qualitatively identical to the non-detuned case, the $g^{(2)}_{\parallel}$ data exhibits symmetrical oscillations around $g^{(2)}_{HOM}=1$. The oscillation amplitude decreases as the delay enlarges, and its lower envelope closely resembles to the non-detuned trace. Under a magnified view (Fig. 2d), the oscillation is discerned to have a period of $1/\Delta$, similarly to the asynchronous TPI result between frequency-offset laser fields 19. As a display of the HOM effect, photon coalescence leads $g^{(2)}_{\parallel}(0)$ to be substantially lower than the oscillation minima. Figure 2f shows the TPI visibility, whose trace has a substantially different appearance from that in Fig. 2e. The $\mathcal{V}$ value oscillates (negative data not shown) with the oscillation maxima upper-bounded to the TPI baseline measured in the non-detuning case. Nevertheless, the $\mathcal{V}$ maximum at $\tau=0$ is about identical to that of the non-detuned case.
@@
To explain the above observations, we resort to the recently introduced RF model 16 that uniquely treats all the RF photons as a result of absorption and re-emission. Considering the TLE and its photon emission together, this RF model adopts a pure-state description for the joint system of the TLE and its emission, although the TLE or the RF emission separately is well known to be in a mixed state. As derived in Appendix A, the quantum state of the light-matter system can be written as
|$\ket{\psi}=\sqrt{p_{0}}\ket{0,g}+\sqrt{p_{1}}{\color[rgb]{0,0,0}e^{-i2\pi\nu t}}\frac{\ket{0,e}+\ket{1,g}}{\sqrt{2}}$ (2)
where the state |$\ket{0}$ (|$\ket{1}$) corresponds to the absence (presence) of a freshly emitted photon, |$\ket{g}$ (|$\ket{e}$) the TLE’s ground (excited) state and $\nu$ the driving laser’s frequency that can be detuned from the TLE’s transition frequency $\nu_{0}$. The terms $p_{0}$ and $p_{1}$ ($p_{0}+p_{1}=1$) represent the probabilities of the TLE system occupying the system ground state |$\ket{0,g}$ and the joint excited state (|$\left(\ket{0,e}+\ket{1,g}\right)/\sqrt{2}$, respectively. The latter is in a superposition form conveying the insight that at any given instant of time, it is not possible to know whether the TLE stays in its excited state |$\ket{e}$ or has decayed to the ground state |$\ket{g}$ and emitted a photon 16. Naturally, |$\ket{1}$ represents a photon emitted by the TLE, implying that the property of the photons is independent of excitation detuning. In the absence of pure dephasing, we have
$p_{0}=\frac{1}{\left(1+2\eta_{ab}\bar{n}\right)},$ (3)
where $\eta_{ab}$ is the TLE’s absorption efficiency at weak pump limit and depends on the excitation detuning.
Eq. 2 allows us to infer the RF’s first-order coherence to be $g^{(1)}(\tau)=p_{0}e^{-i2\pi\nu\tau}$ for $\tau\gg T_{1}$, via evaluation of the interference outcome between the pure quantum state with a delayed copy of itself 16. Through Fourier transform of $g^{(1)}(\tau)$, we readily obtain that the detuned RF spectrum contains a laser-like component, with a spectral weight of $p_{0}$, that inherits the driving laser’s frequency ($\nu$) and coherence time ($T_{L}$). At weak pump limit, the fraction of the laser-like component can be arbitrarily close to unity. Hence, the vast RF spectral distinguishability shown in Fig. 1b can be attributed to the interference of the phase-locked temporal modes of single photons with properties identically defined by the QD.
@@
Fig. 3: Visualization of two-photon coincidence contributions. a, Two mutually-detuned RF beams, with global phases of $\phi_{a}$ and $\phi_{b}$, respectively, meet at a 50/50 beam splitter, and their interference is recorded by two single-photon detectors; b, Four possible cases involving two input photons may contribute to a coincidence event with the two detectors clicking respectively at time $t_{1}$ and $t_{2}=t_{1}+\tau$. Cases (ii) and (iii), in which each input arm contributes one photon, share the common global phase ($\phi_{a}+\phi_{b}$) and therefore their quantum probability amplitudes interfere. Cases that involve three or more input photons exist and are not shown.
@@
The dual-color HOM data can be elegantly explained through the schematic shown in Fig. 3. Without loss of generality, we fix the entry of the red (blue) detuned RF signal to the top (bottom) input of the 50/50 beam splitter. Their global phases are respectively depicted as $\phi_{a}$ and $\phi_{b}$, and are not mutually locked. A coincidence of photon clicks at $t_{1}$ and $t_{2}=t_{1}+\tau$ by two detectors can arise from four possible cases: both photons came from the same entry (i and iv) or they entered from different entries (ii and iii). Here, we temporally ignore cases involving three or more input photons to avoid complication. Among the four possible cases, cases (ii) and (iii) share the same global phase ($\phi_{a}+\phi_{b}$) and therefore interference occurs after their transformation through the beam splitter. The phase difference is $2\pi\Delta\tau+\pi$, where $2\pi\Delta\tau$ arises from the RF detuning and an extra $\pi$ phase gains from the beam splitter. The interference will cause the total coincidence probability to oscillate around an intermediate value that is defined by cases (i) and (iv).
@@
where $\mathcal{M}=|\langle 1_{+}|1_{-}\rangle|^{2}$ represents indistinguishability between temporally-aligned single photon states (|$\ket{1_{\pm}}$) generated under opposite detunings. We remark that states |$\ket{1_{\pm}}$ represent spontaneous emission and therefore their indistinguishability is affected by the QD environment fluctuation but independent of the excitation detuning. $\mathcal{M}$ can also be adjusted through the half-wave plate in the experimental setup (Fig. 1). With orthogonal polarizations, $\mathcal{M}=0$ and $g^{(2)}_{\perp}(0)$ has a theoretical minimum of 0.5.
@@
The first term $\frac{1}{2}$ arises from cases (i) and (iv), i.e., both photons within a coincidence event came from the same beam splitter input. The second term represents the interference between the quantum probability amplitudes of cases (ii) and (iii), but which can no longer cancel completely because of the non-negligible probability of each input arm contributing two photons. The detailed derivation of Eq. 5 is presented in Appendix B.
@@
We simulate the HOM experiments using Eq. 6, with results shown as solid lines in Fig. 2. The auto-correlation function is directly taken from the experimental HBT data using $g^{(2)}(\tau)=\sqrt{g_{+}^{(2)}(\tau)g_{-}^{(2)}(\tau)}$, where $g_{\pm}^{(2)}(\tau)$ are the measured auto-correlation functions under the corresponding positive ($+\frac{\Delta}{2}$) or negative ($-\frac{\Delta}{2}$) excitation detuning. We use $\mathcal{M}=0.94$ ($\mathcal{M}=0$) for the parallel (orthogonal) polarization setting. We also take into account the finite coherence time between the driving lasers, which deteriorates the TPI visibility at large delays. As reported in Figs. 2ad, the simulation faithfully reproduces all the measurement data for both polarization settings, both excitation detunings and all delays.
@@
In the above simulation, we choose $p_{0}=0.92$ corresponding to an excitation flux of $\bar{n}=0.05$. As $p_{0}$ decreases with a strengthening driving intensity, we expect the oscillation amplitude of the detuned HOM correlation to dampen according to Eq. 5. To test, we measure power-dependent dual-color HOM traces with the results (symbols) shown in Fig. 4a. Fitting the correlation oscillations using Eq. 5 (solid lines), we extract the oscillation amplitude ($\frac{1}{2}\mathcal{M}p_{0}^{2}$) and plot its dependence on the excitation flux in Fig. 4b. The measured dependence (symbol) is in agreement with the theoretical simulation (dashed line) using Eq. 3.
Fig. 5 reports the measured 0-delay HOM correlation values as a function of excitation detuning. We observe a gradual increase of $g^{(2)}_{\parallel}(0)$, which is expected from the RF’s deteriorating single-photon purity at large excitation detunings (see Fig. 1b). However, $g^{(2)}_{\perp}(0)$ exhibits an opposite trend, and unexpectedly decreases as the excitation detuning exceeds 0.5 GHz. At $\Delta=4$ GHz, $g^{(2)}_{\perp}(0)$ is measured to be just 0.31, which is significantly below the theoretical minimum of 0.5 for distinguishable photons. Often, $g^{(2)}_{\perp}(0)<0.5$ can be caused by an intensity imbalanced AMZI, which is a common flaw in a HOM setup 15. For example, at the extreme condition with one AMZI arm completely blocked, the HOM interferometer becomes equivalent to a HBT setup. We have ruled out the AMZI as the possible cause for $g^{(2)}_{\perp}(0)<0.5$ in our experiment. Consequently, $g^{(2)}_{\perp}(0)<0.5$ can be caused only by the QD source and is probably related to the large excitation detuning. While the exact mechanism is yet to be pinned down, it is likely to have a similar effect on the parallel polarization setting and hence Eq. 1 can remain meaningful for calculating the TPI visibility.
The experimental TPI visilibity is shown as solid squares in Fig. 5. When $\Delta>0.5$ GHz, $\mathcal{V}(0)$ decreases monotonically with the excitation detuning. The TPI visibility $\mathcal{V}(0)$ is closely linked to photon indistinguishability. In an ideal scenario of $g^{(2)}(0)=0$, it is straightforward to deduce $\mathcal{V}(0)=\mathcal{M}$ from Eqs. 1 and 6. As all RF photons are treated as spontaneous emission, the pure-state RF model (Eq. 2) immediately predicts their indistinguishability to be independent of excitation detuning. Apparently, the measured $\mathcal{V}(0)$ does not follow the predicted behavior for photon distinguishability. A number of factors can contribute to this discrepancy. Firstly, single-photon purity degrades at large detunings due to the increasing laser background. Secondly, the detectors’ time resolution prevents from resolving the HOM interference dip, especially when the dip is modulated by a large detuning frequency. Finally, the measured $g^{(2)}_{\perp}(0)<0.5$ at large detunings effectively reduce the measured $\mathcal{V}(0)$ value. By taking into account these factors, we construct a model (see Appendix D) and calculate the post-selective TPI visibility with results as the red line in Fig. 5. Here, each required $g^{(2)}(0)$ value is interpolated from the experimental measurement shown in Fig. 1b. In accordance with the RF model (Eq. 2), we use a fixed value of $\mathcal{M}=0.94$, which was conservatively chosen based on the first-order coherence measurement shown in Appendix C. The result (green line) is in excellent agreement with the experimental data (squares), thereby supporting the RF model based on the pure-state description.
@@
Fig. 4: Excitation flux dependence of the HOM correlation under the parallel polarization setting. The excitation detuning is fixed at $\Delta=0.5$ GHz. a, HOM correlation traces. Symbols (solid lines) denote experimental data (theoretical fits). b, The oscillation amplitude (symbols) extracted from the data shown in panel a as well as the fit (dashed line) using Eq. 3.

### v3

@@ Abstract
The radiative linewidth of a two-level emitter (TLE) fundamentally limits the bandwidth available for quantum information processing. Despite its importance, no prior experiment has systematically examined how driving detuning affects the indistinguishability of photons scattered from a TLE—a parameter critical for photonic quantum computing. Here, we perform post-selective two-photon interference measurements between mutually detuned resonance fluorescence signals from an InAs quantum dot embedded in a micropillar cavity. At small mutual laser detunings ($\leq 0.5$ GHz), the results are accurately described by the pure-state model [Nat. Commun. 16, 6453 (2025)], which treats all resonance-fluorescence photons as spontaneous emission. At larger detunings, we uncover an anomalous feature in the two-photon interference, where the normalised second-order correlation function under orthogonal polarisations yields $g^{(2)}_{\perp}(0)<0.5$.
@@
Impinging upon a semi-transparent beam-splitter, two indistinguishable single photons will coalesce into either of the two output ports, and never exit separately. Referred to as Hong-Ou-Mandel (HOM) effect 1, this phenomenon is the foundation for quantum information protocols such as repeater-assisted quantum communication 2 and linear-optic quantum computing 3. Single photons can be generated through exciting an atom-like system 4, 5, 6, which by nature emits one photon at a time. In this regard, semiconductor quantum dots (QDs) are arguably the most attractive platform because they can be integrated into monolithic structures for cavity enhancement 5, 7, 8, 9, 10 and highly indistinguishable single photons 11, 12.
@@
Resonant excitation into the lowest excited state allows a QD system to be simplified as a two-level emitter (TLE) 13. Under a weak coherent pump (Heitler regime), the resulting resonance fluorescence (RF) signal locks to the driving laser’s coherence 14, 15, 16, 17 while exhibiting single-photon characteristics. To a first approximation, the QD appears to behave like a passive and elastic scatterer, which led to an optimistic idea to circumvent environment-induced dephasing inherent in all solid-state systems and to generate single photons with indistinguishability “fundamentally” defined by the excitation laser 15. A question then arises. In the presence of an excitation detuning, will the single photons remain indistinguishable from their counterparts generated under strictly resonant excitation? An affirmative answer would be favorable for quantum information processing because it will then permit arbitrary modulation of the driving laser and the shaping of the RF beam accordingly without loss of photon indistinguishability. However, no experiment has so far been dedicated to explicitly examining this question, while substantial efforts have been made to verify photon indistinguishability between remote quantum emitters 18, 19, 20, 21, 22.
@@
Fig. 1: Dual-color two-photon interference (TPI) setup. The RF signals are collected in the same polarization as the excitation lasers, and no spectral filtering is applied to them in any of the TPI experiments. a, High-resolution RF spectra measured with a Fabry-Pèrot interferometer (FPI); The dashed line represents the QD-micropillar reflectance spectrum measured under $\bar{n}=0.05$. b, The auto-correlation function $g^{(2)}(0)$ values measured under $\bar{n}=0.05$ using a Hanbury Brown-Twiss setup (HBT). The red dashed line shows the cavity reflectivity under a strong excitation of 465 nW, corresponding to $\bar{n}=139$. BS: beam splitter; SPD: single-photon detector.
@@
Here, we subject an InAs QD-micropillar device to dual-color excitation, i.e., from two different wavelengths, using the setup shown in Fig. 1, and evaluate indistinguishability of mutually-detuned RF signals using post-selective two-photon interference (TPI) 23. Two continuous-wave (CW) lasers (not shown) are each modulated into square pulses of 595 ns duration, separated by 595 ns off periods, and are then temporally interleaved to sequentially excite the QD device. Symmetrical detuning is applied: one laser is detuned by $\frac{\Delta}{2}$ while the other by $\frac{\Delta}{2}$ from the QD resonance ($\nu_{0}$). The resulting RF tracks the lasers’ frequencies and consists of RF signal pulses of alternating colors. Without applying any spectral filtering, the RF signals enter an asymmetric Mach-Zehnder interferometer (AMZI) of 595 ns delay that matches the laser carving period, so that the RF signals of different colors are temporally aligned when meeting at the AMZI’s exit beam splitter. The interference results are registered by a pair of superconducting nanowire single photon detectors (SPDs) of 59 ps temporal resolution.
[…]
Further experimental details and measurement data, including the pulse-carving of the excitation laser as well as the determinations of the QD’s radiative lifetime ($T_{1}$) and fine-structure splitting, are provided in Appendix B.
@@
Before the dual-color experiment, we characterized the QD under excitation by a single laser. The excitation flux is kept at $\bar{n}=0.05$, which is conservatively set low to operate in the Heitler regime. Fig. 1a shows a set of high-resolution RF spectra measured by a Fabry-Pèrot interferometer (FPI). At each detuning, the RF peak inherits the laser’s optical frequency and has a full-width at half-maximum width (FWHM) that is limited by the FPI resolution ($20$ MHz). The peak intensity decreases as the detuning widens, which is a result of the reduced cavity enhancement. At the measurement step-size of 0.5 GHz, the adjacent RF spectra are fully distinguished. Fig. 1b shows the corresponding auto-correlation function results, measured with a Hanbury Brown-Twiss (HBT) setup. Across the detuning range, $g^{(2)}(0)$ is consistently lower than 0.5, indicating the non-classical nature of each RF signal. The lowest $g^{(2)}(0)=0.026$ is obtained at 0-detuning. It gradually deteriorates as the detuning increases, and exceeds $0.10$ when the detuning reaches $\pm 2$ GHz. We attribute the deterioration in single-photon purity to the increased laser background, as evidenced by the cavity reflectivity (red dashed line), which was measured under strong laser excitation.
@@
We use the setup in Fig. 1 for dual-color excitation experiments. By rotating the half-wave plate ($\lambda/2$) in the long arm of the AMZI, the polarization alignment between the incoming photon beams can be made either parallel or orthogonal. In the latter case, photons are fully distinguishable. We measure the normalized second-order correlation functions under both polarization settings, from which we calculate the TPI visibility ($\mathcal{V}$) using 23
@@
As a control experiment, we set the two lasers to be nominally identical in wavelength, i.e., $\Delta=0$ GHz, and Fig. 2a shows the measured HOM traces (dots). Here, the coincidence probabilities are normalized to the average rate at long delays ($\tau>200$ ns). With orthogonal polarizations, the photons are fully distinguishable and therefore no interference takes place. Hence, the corresponding $g^{(2)}_{\perp}(\tau)$ trace displays a single, sharp dip with a depth of $\sim$0.5 and a width of about 100 ps, arising from the single-photon nature of the RF streams. When the polarizations are aligned, photons become largely indistinguishable, two-photon coalescence takes place thus resulting in a more pronounced HOM dip at $\tau=0$. Additionally, the correlation function exhibits a broad suppression extending over a timescale of $\pm 200$ ns, corresponding to the mutual coherence time of the two pump lasers. Over nanosecond timescale (Fig. 2c), $g_{\perp}^{(2)}$ and $g_{\parallel}^{(2)}$ exhibit flat values for $\tau\gg T_{1}$ but with $g_{\parallel}^{(2)}$ being considerably lower than $g_{\perp}^{(2)}$. This is a common observation in the HOM measurements for resonantly excited QDs 16. Using Eq. 1, we obtain the TPI visibility shown in Fig. 2e.
@@
Next, we measure the second-order HOM correlation functions under dual-color excitation, with the result of a $\Delta=0.5$ GHz detuning shown in Fig. 2b. While the $g^{(2)}_{\perp}$ data is qualitatively identical to the non-detuned case, the $g^{(2)}_{\parallel}$ data exhibits symmetrical oscillations around $g^{(2)}_{HOM}=1$. The oscillation amplitude decreases as the delay enlarges, and its lower envelope closely resembles the non-detuned trace. Under a magnified view (Fig. 2d), the oscillation is discerned to have a period of $1/\Delta$, similarly to the asynchronous TPI result between frequency-offset laser fields 25. As a display of the HOM effect, photon coalescence leads $g^{(2)}_{\parallel}(0)$ to be substantially lower than the oscillation minima. Figure 2f shows the TPI visibility, whose trace has a substantially different appearance from that in Fig. 2e. The $\mathcal{V}$ value oscillates (negative data not shown) with the oscillation maxima upper-bounded to the TPI baseline measured in the non-detuning case. Nevertheless, the $\mathcal{V}$ maximum at $\tau=0$ is about identical to that of the non-detuned case.
@@
To explain the above observations, we resort to the recently introduced RF model 17 that uniquely treats all the RF photons as a result of absorption and re-emission. Considering the TLE and its photon emission together, this RF model adopts a pure-state description for the joint system of the TLE and its emission, although the TLE or the RF emission separately is well known to be in a mixed state. As derived in Appendix A1, the quantum state of the light-matter system can be written as
[…]
Equation 2 provides a unified description of resonance fluorescence, reproducing both the laser-like spectrum and photon antibunching shown in Fig. 1b and c. First, the quantum state can be shown to exhibit photon antibunching, as characterized by the second-order correlation function
$g^{(2)}(\tau)=\frac{\langle a^{\dagger}(t)a^{\dagger}(t+\tau)a(t+\tau)a(t)\rangle}{\langle a^{\dagger}(t)a(t)\rangle^{2}},$
which yields $g^{(2)}(0)=0$ owing to the absence of two-photon probability. Second, we can infer the power spectrum by working out the first-order coherence function,
$g^{(1)}(\tau)=\frac{\langle a^{\dagger}(t)a(t+\tau)\rangle}{\langle a^{\dagger}(t)a(t)\rangle}=p_{0}e^{-i2\pi\nu\tau},$
valid for $T_{1}\ll\tau\ll T_{L}$, where $T_{L}$ denotes the driving laser’s coherence time. The Fourier transform of $g^{(1)}(\tau)$ then yields a power spectrum containing a narrow, laser-like component centered at $\nu$ with a spectral weight of $p_{0}$. Specifically, $p_{0}$ is dependent on the excitation power
$p_{0}=\frac{1}{1+2\eta_{ab}\bar{n}},$ (3)
where $\eta_{ab}$ denotes the TLE’s absorption efficiency in the weak-drive regime.
@@
Fig. 3: Visualization of two-photon coincidence contributions. a, Two mutually-detuned RF beams, with global phases of $\phi_{a}$ and $\phi_{b}$, respectively, meet at a 50/50 beam splitter, and their interference is recorded by two single-photon detectors; b, Four possible cases involving two input photons may contribute to a coincidence event with the two detectors clicking respectively at time $t_{1}$ and $t_{2}=t_{1}+\tau$. Cases (ii) and (iii), in which each input arm contributes one photon, share the common global phase ($\phi_{a}+\phi_{b}$) and therefore their quantum probability amplitudes interfere. For simplicity, cases involving three or more input photons are omitted from the diagram, but they are fully accounted for in Eq. 6.
@@
The dual-color HOM data can be elegantly explained through the schematic shown in Fig. 3. Without loss of generality, we fix the entry of the red (blue) detuned RF signal to the top (bottom) input of the 50/50 beam splitter. Their global phases are respectively depicted as $\phi_{a}$ and $\phi_{b}$, and are not mutually locked. A coincidence of photon clicks at $t_{1}$ and $t_{2}=t_{1}+\tau$ by two detectors can arise from four possible cases: both photons came from the same entry (i and iv) or they entered from different entries (ii and iii). Here, we temporarily ignore cases involving three or more input photons to avoid complications. Among the four possible cases, cases (ii) and (iii) share the same global phase ($\phi_{a}+\phi_{b}$) and therefore interference occurs after their transformation through the beam splitter. The phase difference is $2\pi\Delta\tau+\pi$, where $2\pi\Delta\tau$ arises from the RF detuning and an extra $\pi$ phase gains from the beam splitter. The interference will cause the total coincidence probability to oscillate around an intermediate value that is defined by cases (i) and (iv).
@@
[…]
@@
The first term $\frac{1}{2}$ arises from cases (i) and (iv), i.e., both photons within a coincidence event came from the same beam splitter input. The second term represents the interference between the quantum probability amplitudes of cases (ii) and (iii), but which can no longer cancel completely because of the non-negligible probability of each input arm contributing two photons. The detailed derivation of Eq. 5 is presented in Appendix A2.
@@
We simulate the HOM experiments using Eq. 6, with results shown as solid lines in Fig. 2. The auto-correlation function is directly taken from the experimental HBT data using $g^{(2)}(\tau)=\sqrt{g_{+}^{(2)}(\tau)g_{-}^{(2)}(\tau)}$, where $g_{\pm}^{(2)}(\tau)$ are the measured auto-correlation functions under the corresponding positive ($\frac{\Delta}{2}$) or negative ($\frac{\Delta}{2}$) excitation detuning. We use $\mathcal{M}=0.94$ ($\mathcal{M}=0$) for the parallel (orthogonal) polarization setting. We also take into account the finite coherence time between the driving lasers, which deteriorates the TPI visibility at large delays. As reported in Figs. 2ad, the simulation faithfully reproduces all the measurement data for both polarization settings, both excitation detunings and all delays.
@@
In the above simulation, we choose $p_{0}=0.920\pm 0.001$ corresponding to an excitation flux of $\bar{n}=0.050\pm 0.002$. As $p_{0}$ decreases with a strengthening driving intensity, we expect the oscillation amplitude of the detuned HOM correlation to dampen according to Eq. 5. To test, we measure power-dependent dual-color HOM traces with the results (symbols) shown in Fig. 4a. Fitting the correlation oscillations using Eq. 5 (solid lines), we extract the oscillation amplitude ($\frac{1}{2}\mathcal{M}p_{0}^{2}$) and plot its dependence on the excitation flux in Fig. 4b. The measured dependence (symbol) is in agreement with the theoretical simulation (dashed line) using Eq. 3. Figure 4b also plots the HOM dip, $g^{(2)}_{\parallel}(0)$, which degrades monotonically with increasing excitation power but at a slower rate than the oscillation amplitude. The degradation arises from the single-photon purity, deteriorated by the increasing laser background at high pump power 17.
@@
Fig. 4: Excitation flux dependence of the HOM correlation under the parallel polarization setting. The excitation detuning is fixed at $\Delta=0.5$ GHz. a, HOM correlation traces. Symbols (solid lines) denote experimental data (theoretical fits). b, The oscillation amplitudes (solid dots) and the zero-delay HOM dip values (open squares), extracted from the data shown in panel a, along with the fit (dashed line) to the former using Eq. 3.
Fig. 5 (bottom panel) reports the measured 0-delay HOM correlation values as a function of excitation detuning. We observe a gradual increase of $g^{(2)}_{\parallel}(0)$, as expected from the RF’s deteriorating single-photon purity at large excitation detunings (see Fig. 1b). In contrast, $g^{(2)}_{\perp}(0)$ exhibits an opposite trend, and anomalously decreases as the excitation detuning exceeds 0.5 GHz. At $\Delta=4$ GHz, $g^{(2)}_{\perp}(0)$ is measured to be just 0.31—well below the theoretical minimum of 0.5 for fully distinguishable photons. Often, $g^{(2)}_{\perp}(0)<0.5$ can be caused by an intensity imbalanced AMZI, which is a common flaw in a HOM setup 16. As an extreme example, the HOM interferometer reduces to a HBT setup when one AMZI arm is completely blocked. However, we have ruled out AMZI imbalance as the cause of the observed $g^{(2)}_{\perp}(0)<0.5$. Instead, we attribute this anomalous behavior to a possible interplay between the large excitation detuning and the QD’s fine-structure splitting. Although the exact mechanism remains to be fully understood, it likely influences both polarization configurations in a similar manner. Consequently, Eq. 1 remains appropriate for evaluating the two-photon interference (TPI) visibility.
The experimental TPI visibility is shown as solid squares in Fig. 5 (top panel). When $\Delta>0.5$ GHz, $\mathcal{V}(0)$ decreases monotonically with the excitation detuning. The TPI visibility $\mathcal{V}(0)$ is closely linked to photon indistinguishability. In an ideal scenario of $g^{(2)}(0)=0$, it is straightforward to deduce $\mathcal{V}(0)=\mathcal{M}$ from Eqs. 1 and 6. As all RF photons are treated as spontaneous emission, the pure-state RF model (Eq. 2) immediately predicts their indistinguishability to be independent of excitation detuning. Apparently, the measured $\mathcal{V}(0)$ does not follow the predicted behavior for photon distinguishability. A number of factors can contribute to this discrepancy. Firstly, single-photon purity degrades at large detunings due to the increasing laser background. Secondly, the detectors’ time resolution prevents resolving the HOM interference dip, especially when the dip is modulated by a large detuning frequency. Finally, the measured $g^{(2)}_{\perp}(0)<0.5$ at large detunings effectively reduce the measured $\mathcal{V}(0)$ value. By taking into account these factors, we construct a model (see Appendix A3) and calculate the post-selective TPI visibility with results as the red line in Fig. 5. Here, each required $g^{(2)}(0)$ value is interpolated from the experimental measurement shown in Fig. 1b. In accordance with the RF model (Eq. 2), we use a fixed value of $\mathcal{M}=0.94$, which was conservatively chosen based on the first-order coherence measurement shown in Appendix B5. The result (green line) is in excellent agreement with the experimental data (squares), thereby supporting the RF model based on the pure-state description.